# Public-private funding models in open source software development: A case study on scikit-learn

### v2

@@ Abstract
Governments are increasingly allocating funding for open source software (OSS) development to address concerns related to software security, digital sovereignty, and national competitiveness in science and innovation, amongst others. While announcements of governmental funding are generally well-received by OSS developers, we still have a limited understanding of OSS developers evaluate the relative benefits and drawbacks of such funding compared to other types of funding. This paper explores this question through a case study on scikit-learn, a Python library for machine learning, whose funding model combines research grants, commercial sponsorship, community donations, and a €32 million grant from the France’s artificial intelligence strategy. Through 25 interviews with scikit-learn’s maintainers and funders, this study makes two key contributions to research and practice. First, the study illustrates how the maintainers have weaved public and private funding into their project to ensure the continued provision of scikit-learn as a digital public good, as well as the importance of diversified funding and governance protocols for funding to safeguard the community ethos of the project. Second, it offers practical recommendations to various stakeholders. For OSS developer communities, it illustrates the benefits of a diversified funding model in balancing the merits and drawbacks of different funding sources. For companies, it serves as a reminder that sponsoring developers or OSS projects can significantly support OSS maintainers, who often struggle with limited resources and towering workloads. For governments, it emphasises the importance of funding the maintenance of existing OSS in addition to or exclusively funding the development of new OSS libraries or features. The paper concludes with suggestions for future research directions.
@@ 1 Introduction
The increasing interest and involvement of governments in funding OSS development has been generally well-received by OSS developers. However, the question of how OSS developers view the relative benefits and drawbacks of governmental funding compared to other sources of funding remains unexplored. This paper addresses this timely issue through a case study of scikit-learn, a popular Python library for machine learning (ML), which is described as “the Swiss army knife of ML” due to its widespread use in research and industry [9]. scikit-learn is developed by a team of core maintainers, who are primarily based at Inria, the French Institute for Research in Computer Science and Automation, and a global community of volunteers. Its sustainability as a community-led OSS project in the industry-dominated field of artificial intelligence (AI) [10] is in part attributed to its mixed funding model, including research grants, commercial sponsorship, and community donations, as well as the institutional backing of Inria since 2010, which have provided the means to employ full-time maintainers, to provide them with office space and IT resources, and to organise community events, amongst others. November 2021 was a turning point for scikit-learn. Recognising the significance of scikit-learn and OSS in general in AI research and development (R&D), the French government announced a €32 million grant in its AI strategy to support the expansion of scikit-learn into a comprehensive suite of open source tools for ML and AI for the benefit of researchers and developers in France and beyond [11].
@@
Through 25 semi-structured interviews with the scikit-learn maintainers and their public and private funders, this paper explores the historical role of mixed public-private funding in supporting the scikit-learn project. The paper makes two key contributions with implications for research and practice. First, the study contributes novel findings about the design and implementation of a public-private funding model in an OSS project, which remains an unexplored topic in the literature. The analysis sheds light on the respective roles that public and private funders have played in supporting scikit-learn and how its maintainers have balanced and when necessary negotiated their funders’ diverse interests. Second, the study provides practical insights for practitioners. For OSS developer communities, it shows that a diversified funding model can act as an effective means for balancing the merits and drawbacks of different funding sources, as well as for containing the influence of funders, whether they are a company or a public body. For companies, it is a timely reminder that by sponsoring developers or an OSS project, they can make a substantial difference to OSS projects and maintainers, who struggle with towering workloads. For governments, it underscores the importance of funding maintenance of existing OSS and refraining from the novelty bias of funding new OSS libraries or features.
@@ 4.1 Stewards of the community: The incentives and stakes of the maintainers
1
China (23.7%)
USA (33.7%)
USA (34.0%)
2
USA (21.8%)
Germany (7.5%)
India (8.9%)
3
India (7.9%)
UK (7.0%)
Germany (8.5%)
4
Germany (4.4%)
India (6.8%)
France (6.9%)
5
Brazil (3.8%)
France (5.5%)
UK (6.4%)
6
UK (3.7%)
China (5.4%)
Canada (4.5%)
7
Canada (3.3%)
Canada (4.5%)
China (2.6%)
8
France (3.2%)
Netherlands (2.5%)
Japan (2.5%)
9
Japan (2.5%)
Switzerland (2.2%)
Switzerland (2.3%)
10
Russia (1.3%)
Australia (2.0%)
Netherlands (2.2%)
@@
1
Google (1.5%)
Google (2.5%)
Google (3.7%)
2
Microsoft (1.5%)
Microsoft (2.0%)
Microsoft (2.5%)
3
Tencent (0.9%)
Amazon (0.7%)
Inria (0.9%)
4
Alibaba (0.8%)
DeepMind (0.7%)
Meta (0.8%)
5
Amazon (0.8%)
IBM (0.6%)
Amazon (0.8%)
6
ByteDance (0.8%)
Johns Hopkins (0.6%)
Dataiku (0.7%)
7
Baidu (0.6%)
Tencent (0.5%)
UC Berkeley (0.6%)
8
Meta (0.5%)
ETH Zurich (0.5%)
Shopify (0.6%)
9
AWS (0.4%)
Uni. of Washington (0.4%)
MIT (0.6%)
10
Tsinghua University (0.4%)
NeuroData (0.4%)
AWS (0.6%)
@@ 4.2 From Inria to France’s AI strategy: The role of public funding in scikit-learn
The French government’s announcement that it would invest €32 million in the expansion of scikit-learn marked a new phase in the public sector’s involvement in the project. The grant is part of the strategy’s aim to cement France’s national competitiveness in AI and to increase AI adoption in the French economy. The strategy highlights that scikit-learn is an asset, which is used by millions of data scientists on a daily basis and which has been supported by Inria since its beginning. A government official explained that the funding for scikit-learn and related open source tools was intended to increase the adoption of AI throughout the country. By virtue of open source, it was acknowledged that researchers and businesses beyond France would benefit, too. In addition to economic goals, the government official explained that digital sovereignty was a key consideration for funding scikit-learn and the digital commons more widely. This was underscored by secondary documents that interviewees shared with me that contextualised the government’s funding for scikit-learn as part of a broader strategy to develop sovereign digital commons in France and Europe. The same government official explained that there is a “real challenge” with the dependence on technologies provided by US Big Tech companies, even including open source frameworks like PyTorch and TensorFlow, in the field of AI and by investing in the digital commons the government aspired to address such dependencies.
@@ 5.1.1 The design and role of public-private OSS funding models
In addition, it highlights the integral role played by the Inria Foundation as the institutional home of the scikit-learn consortium and the scikit-learn project’s governance protocols, which have served as crucial mechanisms for the maintainers to contain the influence of funders and uphold the project’s community ethos. This finding extends prior work on the role of foundations as “boundary organisations” that enable collaboration between OSS developers and for-profit companies with new evidence of the role of these institutions in handling the politics of money as an instrument of influence in community-led OSS projects [50, 24]. Moreover, the diversified funding model not only ensures financial stability for the project but also acts as a safeguard when disagreements arise between maintainers and funders. This diversification protects the project from the risk of over-reliance on the support of a single company.
@@ 6 Conclusion
Funding plays a critical role in supporting the sustainability of community-led OSS projects. This case study on scikit-learn makes two key contributions with implications for research and practice. First, the study contributes novel findings to the academic literature on emerging OSS funding models. It illustrates how the maintainers of scikit-learn have weaved public and private sources of funding into their project in the interest of ensuring the continued provision of scikit-learn as a digital public good, whilst employing governance protocols to safeguard the community ethos of the project. Second, it provides practical recommendations for OSS practitioners. For OSS developer communities, the case of scikit-learn shows that the diversification of funding can act as a useful mechanism for sustaining a community-led OSS project, whilst safeguarding the community ethos of the project and its independence from the influence of single funders. For companies, it is a timely reminder that by sponsoring an OSS developer or funding an OSS project they can make a significant difference to maintainers. For governments, it underscores the critical importance of refraining from the novelty bias of only funding new innovation and funding the maintenance of existing OSS. Overall, this study underscores the effectiveness of diversified funding models as a pragmatic approach to sustaining community-led OSS projects in competitive industries. Ultimately, the future of OSS funding depends on the development of sustainable strategies that align the interests of funders with the needs and norms of the global developer communities that build and maintain our open source digital infrastructure.

### v3

@@ Abstract
Governments are increasingly funding open source software (OSS) development to address concerns regarding software security, digital sovereignty, and national competitiveness in science and innovation. While announcements of governmental funding are generally well-received by OSS developers, we still have a limited understanding of how they evaluate the relative benefits and drawbacks of such funding compared to other types of funding. This paper explores this question through a case study on scikit-learn, a Python library for machine learning, whose funding combines research grants, commercial sponsorship, community donations, and a €32 million grant from France’s artificial intelligence strategy. Through 25 interviews with scikit-learn’s maintainers and funders, this study makes two key contributions to research and practice. First, the study contributes novel findings about the design and implementation of a public-private funding model in an OSS project. It sheds light on the respective roles that public and private funders have played in supporting scikit-learn, and the processes and governance mechanisms employed by the maintainers to balance their funders’ diverse interests and to safeguard community interests. Second, it offers practical recommendations. For OSS developer communities, it illustrates the benefits of a diversified funding model for balancing the merits and drawbacks of different funding sources and mitigating dependence on single funders. For companies, it serves as a reminder that sponsoring developers or OSS projects can significantly help maintainers, who often struggle with limited resources and towering workloads. For governments, it emphasises the importance of funding the maintenance of existing OSS in addition to funding the development of new software or features. The paper concludes with suggestions for future research.
@@ 1 Introduction
The increasing interest and involvement of governments in funding OSS development has been generally well-received by OSS developers. However, the question of how OSS developers view the relative benefits and drawbacks of governmental funding compared to other sources of funding remains unexplored. This paper addresses this timely issue through a case study on scikit-learn, a popular Python library for machine learning (ML), which is described as “the Swiss army knife of ML” due to its widespread use in research and industry [9]. scikit-learn is developed by a team of core maintainers, who are primarily based at Inria, the French Institute for Research in Computer Science and Automation, and a global community of volunteers. Its sustainability as a community-led OSS project in the industry-dominated field of artificial intelligence (AI) [10] is in part attributed to its mixed funding model, including research grants, commercial sponsorship, and community donations, as well as the institutional backing of Inria since 2010, which have provided the means to employ full-time maintainers, to provide them with office space and IT resources, and to organise community events, amongst others. November 2021 was a turning point for scikit-learn. Recognising the significance of scikit-learn and OSS in general in AI research and development (R&D), the French government announced a €32 million grant in its AI strategy to support the expansion of scikit-learn into a comprehensive suite of open source tools for ML and AI for the benefit of researchers and developers in France and beyond [11].
@@
Through 25 semi-structured interviews with the scikit-learn maintainers and their public and private funders, this paper explores the historical role of mixed public-private funding in supporting the scikit-learn project. The paper makes two key contributions with implications for research and practice. First, the study contributes novel empirical findings about the design and implementation of a public-private funding model in an OSS project. The analysis sheds light on the respective roles that public and private funders have played in supporting scikit-learn and how its maintainers have balanced and when necessary negotiated their funders’ diverse interests. Second, the study provides practical insights for practitioners. For OSS developer communities, it shows that a diversified funding model can act as an effective means for balancing the merits and drawbacks of different funding sources, as well as for containing the influence of any single funder, whether they are a company, a government body, or other. For companies, it is a timely reminder that by sponsoring developers or an OSS project, they can make a substantial difference to OSS maintainers, who struggle with towering workloads and limited resources. For governments, it underscores the importance of funding maintenance of existing OSS and refraining from the novelty bias of funding new software or features.
@@ 4.1 Stewards of the community: The incentives and stakes of the maintainers
1 China (23.7%) USA (33.7%) USA (34.0%)
2 USA (21.8%) Germany (7.5%) India (8.9%)
3 India (7.9%) UK (7.0%) Germany (8.5%)
4 Germany (4.4%) India (6.8%) France (6.9%)
5 Brazil (3.8%) France (5.5%) UK (6.4%)
6 UK (3.7%) China (5.4%) Canada (4.5%)
7 Canada (3.3%) Canada (4.5%) China (2.6%)
8 France (3.2%) Netherlands (2.5%) Japan (2.5%)
9 Japan (2.5%) Switzerland (2.2%) Switzerland (2.3%)
10 Russia (1.3%) Australia (2.0%) Netherlands (2.2%)
@@
1 Google (1.5%) Google (2.5%) Google (3.7%)
2 Microsoft (1.5%) Microsoft (2.0%) Microsoft (2.5%)
3 Tencent (0.9%) Amazon (0.7%) Inria (0.9%)
4 Alibaba (0.8%) DeepMind (0.7%) Meta (0.8%)
5 Amazon (0.8%) IBM (0.6%) Amazon (0.8%)
6 ByteDance (0.8%) Johns Hopkins (0.6%) Dataiku (0.7%)
7 Baidu (0.6%) Tencent (0.5%) UC Berkeley (0.6%)
8 Meta (0.5%) ETH Zurich (0.5%) Shopify (0.6%)
9 AWS (0.4%) Uni. of Washington (0.4%) MIT (0.6%)
10 Tsinghua University (0.4%) NeuroData (0.4%) AWS (0.6%)
@@ 4.2 From Inria to France’s AI strategy: The role of public funding in scikit-learn
The French government’s announcement that it would invest €32 million in the expansion of scikit-learn marked a new phase in the public sector’s involvement in the project. The grant is part of the strategy’s aim to cement France’s national competitiveness in AI and to increase AI adoption in the French economy. The strategy highlights that scikit-learn is an asset, which is used by millions of data scientists on a daily basis and which has been supported by Inria since its beginning. A government official explained that the funding for scikit-learn and related open source tools was intended to increase the adoption of AI throughout the country. By virtue of being open source, it was acknowledged that researchers and businesses beyond France would benefit, too. In addition to economic goals, the government official explained that digital sovereignty was a key consideration for funding scikit-learn and the digital commons more widely. This was underscored by secondary documents that interviewees shared with me that contextualised the government’s funding for scikit-learn as part of a broader strategy to develop sovereign digital commons in France and Europe. The same government official explained that there is a “real challenge” with the dependence on technologies provided by US Big Tech companies, even including open source frameworks like PyTorch and TensorFlow, in the field of AI and by investing in the digital commons the government aspired to address such dependencies.
@@ 5.1.1 The design and role of public-private OSS funding models
In addition, it highlights the integral role played by the Inria Foundation as the institutional home of the scikit-learn consortium and the scikit-learn project’s governance protocols, which have served as crucial mechanisms for the maintainers to contain the influence of funders and uphold the project’s community ethos. This finding extends prior work on the role of foundations as “boundary organisations” that enable collaboration between OSS developers and for-profit companies with new evidence of the role of these institutions in handling the politics of money as an instrument of influence in community-led OSS projects [50, 24]. Moreover, the diversified funding model not only ensures financial stability for the project but also acts as a safeguard when disagreements arise between maintainers and funders. This diversification protects the project from the risk of over-reliance on the support of a single funder.
@@ 6 Conclusion
Funding plays a critical role in supporting the sustainability of community-led OSS projects. This case study on scikit-learn makes two key contributions with implications for research and practice. First, the study contributes novel findings on emerging OSS funding models to the academic literature. It illustrates how the maintainers of scikit-learn have weaved public and private sources of funding into their project in the interest of ensuring the continued provision of scikit-learn as a digital public good, whilst employing governance protocols to safeguard the community ethos of the project. Second, it provides practical recommendations for OSS practitioners. For OSS developer communities, it shows that the diversification of funding can act as a useful mechanism for sustaining a community-led OSS project, whilst safeguarding the community ethos of the project and its independence from the overreaching influence of single funders, whether they are a company, a government body, or other. For companies, it is a timely reminder that by sponsoring an OSS developer or funding an OSS project they can make a significant difference to maintainers. For governments, it underscores the critical importance of refraining from the novelty bias of only funding new innovation and funding the maintenance of existing OSS. Overall, this study underscores the effectiveness of diversified funding models as a pragmatic approach to sustaining community-led OSS projects in competitive industries. Ultimately, the future of OSS funding depends on the development of sustainable strategies that align the interests of funders with the needs and norms of the global developer communities that build and maintain our open source digital infrastructure.